OpenStudy (anonymous):

Change the following mixed expressions to fractions and then simplify the complex fractions. (xy) subract (4x^5) / (x^2-xy)

OpenStudy (dumbcow):

\[\frac{xy - 4x^{5}}{x^{2} -xy} = \frac{x(y-4x^{4})}{x(x-y)} = \frac{y-4x^{4}}{x-y} \]